A random variable X has Poisson distribution with mean equal to 0.4. What is the
probability that the random variable is greater than zero
1
Expert's answer
2020-06-01T19:21:43-0400

We have,


X ~ Poisson("\lambda"), "\lambda" > 0


For Poisson distribution mean = "\lambda", then by the problem we have "\lambda" = 0.4.


The p.m.f. of X is given by,


P(X = x) = "\begin{cases}
 \frac{e^{-0.4}.(0.4)^x}{x!} &\text{if } x=0,1,2,...\infty \\
 0 &otherwise
\end{cases}"


The probability that the random variable is greater than 0


= P(X > 0)


= 1 - P(X = 0) [since the total probability is 1]


= 1 - "\frac{e^{-0.4}.(0.4)^0}{0!}"


= 1 - 0.6703 = 0.3297 (rounded to 4 decimal places)


Answer: The probability that the random variable is greater than zero is 0.3297.